OpenStudy (anonymous):

Find an equation of the line that is tangent to the circle x^2 +y^2=25 at the point p(3,4)

OpenStudy (anonymous):

You know derivative?

OpenStudy (anonymous):

Ok lets find the slope of the line from the center of the circle to the point w/c is given. What do u think it would be?

OpenStudy (anonymous):

so it will be (0,0) and (3,4)?

OpenStudy (anonymous):

\[m_{\perp} =\frac{4-0}{3-0}\]

OpenStudy (anonymous):

so the slope will be 4/3 isn't

OpenStudy (anonymous):

It is the slope of the radius to the point (3, 4)

OpenStudy (anonymous):

Now find the slope w/c is perpendicular to the above slope. Tell me what it would be?

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

YEP now u have the slope and the point. what will be the line?

OpenStudy (anonymous):

so we substitute it in the point-slope formula? y-4=-3/4(x-3)

OpenStudy (anonymous):

Great, done.

OpenStudy (anonymous):

but why do we need to make the slope be perpendicular?

OpenStudy (anonymous):

w/c slope?

OpenStudy (anonymous):

oh, ok thankyou very much

OpenStudy (anonymous):

well come